the angle between an incident ray and mirror is 30°. What is the total angle turned by the Ray?


1
Expert's answer
2021-04-29T10:43:11-0400

The total angle between the incident and reflected rays is


"\alpha=180°-2\theta=180°-2·30°=120°."